# High CO/H2 ratios supports an exocometary origin for a CO-rich debris disk

### v1

@@ 1 Introduction
Exocometary belts were initially assumed to be gas-free; however a variety of atomic species (eg. Hobbs et al., 1985) as well as CO have been detected in over 20 exocometary systems (brennan_low_2024; Cataldi et al., 2023; Marino et al., 2020; Kral et al., 2017; Moór et al., 2017; Matrà et al., 2017b; Lieman-Sifry et al., 2016; Kóspál et al., 2013; Troutman et al., 2011; Dent et al., 2005; Zuckerman, 1995, eg.). Thus far CO is the only molecule detected in exocometary gas due to its resistance to photodissociation, chemical stability, and easy detectability (Visser et al., 2009). CO has strong rovibronic transitions, and the typically cold gas present in exocometary belts is detectable in UV rovibronic absorption spectra by the Hubble Space Telescope (HST) (eg. Roberge et al., 2000), rotational transitions in emission detectable in mm with ALMA (eg. Matrà et al., 2017b; Cataldi et al., 2023; Rebollido et al., 2022) and IR absorption from rovibrational transitions as detected using the Gemini South Telescope and the NASA Infrared Telescope Facility around $\beta$ Pictoris (Troutman et al., 2011).
@@
We use the newly upgraded, high-resolution CRIRES+ instrument on the VLT to observe the edge-on, CO-rich disks around HD 110058 and HD 131488. With a spectral resolution $R\sim 100000$ (Leibundgut et al., 2022), CRIRES+ is ideal for detecting line of sight absorption due to CO and H2. Our observations cover multiple rovibrational lines of 12CO from the v=2-0 level. Absorption from CO in edge-on debris disks has been successfully observed by Troutman et al. (2011), who reported low-$J$ CO absorption from the $\beta$ Pictoris disk using IRTF/CSHELL. Worthen et al. (2024) reported a non-detection of CO in absorption toward the edge-on HD 32297 disk using IRTF/iSHELL which was used to estimate the scale height of the disk and place an upper limit on the line-of-sight CO column density. The two disks in this study were measured in absorption in the UV using Hubble by brennan_low_2024. Our CRIRES+ observations allow for CO column density and temperature estimates for both HD 110058 and HD 131488.
In this work, we aim to directly probe the column density of H2 within CO-rich exocometary belts for the first time, and in doing so establish the CO origin. We expect any exocometary $\text{H}_{2}$ to be co-located with and similar temperature as the CO detected through UV absorption spectroscopy in both disks by brennan_low_2024 and in IR in this work. However it is worth noting that H2 should have a larger vertical extent than CO due to CO being more readily photodissociated at the disk’s surface and it’s lower mean molecular weight (Marino et al., 2020; Hughes et al., 2017). Both H2 and CO should be cold and produce narrow absorption lines from their ground vibrational energy levels. We focus on absorption spectroscopy in edge-on belts to use the star as a bright background continuum, and maximise the column density of gas along the line of sight to it. We identified the H2 v=1-0 S(0) line at 2223 nm as the most promising transition probing rovibrational absorption from the ground level of the H2 molecule. This is because, assuming the H2 gas in these systems has a similar temperature to the CO and is in LTE, this line would be the strongest IR transition.
Stellar and disk properties for HD 110058 and HD 131488 are detailed in Table 1. Both are A-type stars (the most common spectral type for CO-rich belts, Moór et al., 2020) and thus provide bright stellar continua against which we searched for absorption. The disks in these systems are also very close to edge-on with inclinations of $85.5^{+2.5}_{-7.2}$ and $82\pm 3$ respectively (Moór et al., 2017; Hales et al., 2022). We choose these two young ($\sim 15$ Myr), CO-rich, edge-on belts as they have the highest CO column densities detected in absorption to date (brennan_low_2024).
@@ 3 Results and Modelling
For both disks, we detect 12CO absorption at a radial velocity consistent with that observed in the UV for HD 110058 by brennan_low_2024. The radial velocity for HD 131488 differs from the one reported in brennan_low_2024, but it is consistent with the CO radial velocity observed in emission with ALMA by Mac Manamon (in press) as part of the ARKS program. For HD 110058, we detect the v=2-0, J=15-14 to J=1-0 and for HD 131488 we detection the v=2-0, J=8-7 to J=1-0 transitions. More CO is present in HD 110058, and thus the absorption lines are deeper, reflected in the axis scaling of Figures 3 and 4. Additionally, the excitation temperature for HD 110058 is higher, allowing more high-J transitions to be populated.
@@ 3.1 Modelling the Reduced CRIRES+ Spectra
We modelled H2 absorption in the normalised spectra using RADIS (Pannier & Laux, 2019). RADIS relies on the HITRAN line list (Gordon et al., 2022) obtained using the HITRAN Application Programming Interface (Kochanov et al., 2016), and partition functions described in Gamache et al. (2021). Given the non-detection, for H2 we assumed the gas is in LTE and that it can be approximated as having a single column density and temperature along the line of sight to the star. On the other hand, for CO we found LTE not to be a sufficiently good description of the data and we therefore adopted a simplified non-LTE with a different kinetic temperature (setting the intrinsic line width through Doppler broadening) from the excitation temperature (setting the ratio of the populations of the various rotational levels, and therefore the relative strength of our detected lines). This is consistent with the findings and the modelling of UV rovibronic lines in the same systems (brennan_low_2024). Although we cannot determine whether H2 is in LTE without more information about the collider densities, the H2 (v=1-0 S(0)) transition has an Einstein A of $2.524\times 10^{-7}\text{s}^{-1}$ while the CO 12CO $v=2\rightarrow 0$ transitions have Einstein A coefficients on the order of $10^{-1}\text{s}^{-1}$. Meanwhile the collisional rate coefficients found in Thi et al. (2013); Quéméner & Balakrishnan (2009) are of comparable orders of magnitude. This means the H2 transition has a lower critical density and is therefore more likely to be in LTE than the detected CO.
@@ 4.2 H2 + self-shielding of CO: likely insufficient in a primordial scenario
Although CI shielding has not been considered previously in the context of a primordial scenario it could extend the CO lifespan, though it is unclear how important this is in younger protoplanetary disks where complex chemistry beyond simply photodissociation involving both CO and C is expected. What we find is that in the radial direction the CI column density in brennan_low_2024 is sufficient to provide shielding over the system lifetime, but it is unclear whether this is possible along the vertical direction as well, where the column density is unconstrained.
@@ 4.3 Challenges remain for the second-generation model
However, brennan_low_2024 detected both CO and CI in absorption in the same two systems as this study, finding that the observed column densities of CI are much lower than predicted by the exocometary release model of Marino et al. (2020) and do not provide sufficient shielding for CO. A similar conclusion was reached by Cataldi et al. (2023), who also determined for 14 exocometary belts that the model of Marino et al. (2020) over predicts the amount of CI. These two results suggest that either some of the assumptions are incorrect, the second generation model lacks some important physics/chemistry, or that the gas is primordial; our work deems the latter option to be unlikely. In conclusion, while our results provide the first direct evidence for likely insufficient amounts of H2 in CO-rich planetesimal belts to shield CO and a markedly different composition to primordial gas in at least one system, current models of second-generation gas release are not yet able to provide a fully self-consistent explanation, and more modelling work and compositional constraints are needed.

### v2

@@ 1 Introduction
Exocometary belts were initially assumed to be gas-free; however a variety of atomic species (eg. Hobbs et al., 1985) as well as CO have been detected in over 20 exocometary systems (Brennan et al., 2025; Cataldi et al., 2023; Marino et al., 2020; Kral et al., 2017; Moór et al., 2017; Matrà et al., 2017b; Lieman-Sifry et al., 2016; Kóspál et al., 2013; Troutman et al., 2011; Dent et al., 2005; Zuckerman, 1995, eg.). Thus far CO is the only molecule detected in exocometary gas due to its resistance to photodissociation, chemical stability, and easy detectability (Visser et al., 2009). CO has strong rovibronic transitions, and the typically cold gas present in exocometary belts is detectable in UV rovibronic absorption spectra by the Hubble Space Telescope (HST) (eg. Roberge et al., 2000), rotational transitions in emission detectable in mm with ALMA (eg. Matrà et al., 2017b; Cataldi et al., 2023; Rebollido et al., 2022) and IR absorption from rovibrational transitions as detected using the Gemini South Telescope and the NASA Infrared Telescope Facility around $\beta$ Pictoris (Troutman et al., 2011).
@@
We use the newly upgraded, high-resolution CRIRES+ instrument on the VLT to observe the edge-on, CO-rich disks around HD 110058 and HD 131488. With a spectral resolution $R\sim 100000$ (Leibundgut et al., 2022), CRIRES+ is ideal for detecting line of sight absorption due to CO and H2. Our observations cover multiple rovibrational lines of 12CO from the v=2-0 level. Absorption from CO in edge-on debris disks has been successfully observed by Troutman et al. (2011), who reported low-$J$ CO absorption from the $\beta$ Pictoris disk using IRTF/CSHELL. Worthen et al. (2024) reported a non-detection of CO in absorption toward the edge-on HD 32297 disk using IRTF/iSHELL which was used to estimate the scale height of the disk and place an upper limit on the line-of-sight CO column density. The two disks in this study were measured in absorption in the UV using Hubble by Brennan et al. (2025). Our CRIRES+ observations allow for CO column density and temperature estimates for both HD 110058 and HD 131488.
In this work, we aim to directly probe the column density of H2 within CO-rich exocometary belts for the first time, and in doing so establish the CO origin. We expect any exocometary $\text{H}_{2}$ to be co-located with and similar temperature as the CO detected through UV absorption spectroscopy in both disks by Brennan et al. (2025) and in IR in this work. However it is worth noting that H2 should have a larger vertical extent than CO due to CO being more readily photodissociated at the disk’s surface and it’s lower mean molecular weight (Marino et al., 2020; Hughes et al., 2017). Both H2 and CO should be cold and produce narrow absorption lines from their ground vibrational energy levels. We focus on absorption spectroscopy in edge-on belts to use the star as a bright background continuum, and maximise the column density of gas along the line of sight to it. We identified the H2 v=1-0 S(0) line at 2223 nm as the most promising transition probing rovibrational absorption from the ground level of the H2 molecule. This is because, assuming the H2 gas in these systems has a similar temperature to the CO and is in LTE, this line would be the strongest IR transition.
Stellar and disk properties for HD 110058 and HD 131488 are detailed in Table 1. Both are A-type stars (the most common spectral type for CO-rich belts, Moór et al., 2020) and thus provide bright stellar continua against which we searched for absorption. The disks in these systems are also very close to edge-on with inclinations of $85.5^{+2.5}_{-7.2}$ and $82\pm 3$ respectively (Moór et al., 2017; Hales et al., 2022). We choose these two young ($\sim 15$ Myr), CO-rich, edge-on belts as they have the highest CO column densities detected in absorption to date (Brennan et al., 2025).
@@ 3 Results and Modelling
For both disks, we detect 12CO absorption at a radial velocity consistent with that observed in the UV for HD 110058 by Brennan et al. (2025). The radial velocity for HD 131488 differs from the one reported in Brennan et al. (2025), but it is consistent with the CO radial velocity observed in emission with ALMA by Mac Manamon (in press) as part of the ARKS program. For HD 110058, we detect the v=2-0, J=15-14 to J=1-0 and for HD 131488 we detection the v=2-0, J=8-7 to J=1-0 transitions. More CO is present in HD 110058, and thus the absorption lines are deeper, reflected in the axis scaling of Figures 3 and 4. Additionally, the excitation temperature for HD 110058 is higher, allowing more high-J transitions to be populated.
@@ 3.1 Modelling the Reduced CRIRES+ Spectra
We modelled H2 absorption in the normalised spectra using RADIS (Pannier & Laux, 2019). RADIS relies on the HITRAN line list (Gordon et al., 2022) obtained using the HITRAN Application Programming Interface (Kochanov et al., 2016), and partition functions described in Gamache et al. (2021). Given the non-detection, for H2 we assumed the gas is in LTE and that it can be approximated as having a single column density and temperature along the line of sight to the star. On the other hand, for CO we found LTE not to be a sufficiently good description of the data and we therefore adopted a simplified non-LTE with a different kinetic temperature (setting the intrinsic line width through Doppler broadening) from the excitation temperature (setting the ratio of the populations of the various rotational levels, and therefore the relative strength of our detected lines). This is consistent with the findings and the modelling of UV rovibronic lines in the same systems (Brennan et al., 2025). Although we cannot determine whether H2 is in LTE without more information about the collider densities, the H2 (v=1-0 S(0)) transition has an Einstein A of $2.524\times 10^{-7}\text{s}^{-1}$ while the CO 12CO $v=2\rightarrow 0$ transitions have Einstein A coefficients on the order of $10^{-1}\text{s}^{-1}$. Meanwhile the collisional rate coefficients found in Thi et al. (2013); Quéméner & Balakrishnan (2009) are of comparable orders of magnitude. This means the H2 transition has a lower critical density and is therefore more likely to be in LTE than the detected CO.
@@ 4.2 H2 + self-shielding of CO: likely insufficient in a primordial scenario
Although CI shielding has not been considered previously in the context of a primordial scenario it could extend the CO lifespan, though it is unclear how important this is in younger protoplanetary disks where complex chemistry beyond simply photodissociation involving both CO and C is expected. What we find is that in the radial direction the CI column density in Brennan et al. (2025) is sufficient to provide shielding over the system lifetime, but it is unclear whether this is possible along the vertical direction as well, where the column density is unconstrained.
@@ 4.3 Challenges remain for the second-generation model
However, Brennan et al. (2025) detected both CO and CI in absorption in the same two systems as this study, finding that the observed column densities of CI are much lower than predicted by the exocometary release model of Marino et al. (2020) and do not provide sufficient shielding for CO. A similar conclusion was reached by Cataldi et al. (2023), who also determined for 14 exocometary belts that the model of Marino et al. (2020) over predicts the amount of CI. These two results suggest that either some of the assumptions are incorrect, the second generation model lacks some important physics/chemistry, or that the gas is primordial; our work deems the latter option to be unlikely. In conclusion, while our results provide the first direct evidence for likely insufficient amounts of H2 in CO-rich planetesimal belts to shield CO and a markedly different composition to primordial gas in at least one system, current models of second-generation gas release are not yet able to provide a fully self-consistent explanation, and more modelling work and compositional constraints are needed.